1.The “per-person” versions of production functions: Write each production functiongiven below in terms of output per persony≡Y /Land capital per personk≡K/L.Showwhat these “per person” versions look like in a graph withkon the horizontal axis andyonthe vertical axis. (Assume ̄A is some fixed positive number.)(a)Y=K1/3L2/3andY=K3/4L1/4(on the same graph)

(b)Y=K

(c)Y=K+ ̄AL

(d)Y=K− ̄AL2.
Answer #1

Divido both sides by 2 y Kill - y = (*)/3 =) y = kl where, y = Y/L = outfart par person. R = k/L = b capital por person. Y =6y=k Divedo both side by L y ly=k y=k+ ĀL Divide both sides by L ERA ly= R + Ā @ Y = K - ALL Divide buth side by L Y=R- 1 you